From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: September 24, 2009

Participants: Daniel, Bill, Simon, Fernando, Beni, Eugene, Brian, Glenn, Mark S., Mark I.

Next Meeting: October 1, 2009 @ 1:30 p.m.

Manpower Update

 - Manpower will be thinning out soon. Stephen Burnett has gone back to
   Hall B. Brian Kross will be off the project for the month of October. 
   We will have to rally available manpower to finish (and start) stack 
   construction. Stay tuned.

General Construction Updates

 - We are presently gearing up to begin stack construction for the
   full-scale prototype. Stay tuned for news and the associated press
   release that will inevitably follow. As part of stack construction, we
   will do a 3D voltage check with a cathode-wire-cathode configuration
   in an Ar-CO2 gas environment to nominal voltage to be sure that there
   are no breakdown problems.
 - Bill has been working on the FDC stack assembly document. He has made
   a draft availale at: /u/group/halld/Engineering/PRELIMINARY DOCUMENTS.
   Folks should look it over and provide feedback. The document will be
   fleshed out during the assembly procedure for the full-scale prototype.
 - We will proceed with regular solder for the rigid-flex soldeing for now.
   Brian will follow up to see if we can get a sample of Indium solder paste.
 - This past week we sealed a single wire board in an Ar/CO2 gas environment
   similar to the configuration of the real FDC detector. Several nominal
   size o-rings were made and used and it turned out to be a big problem to
   keep them in their grooves. Several lessons were learned about how to
   make this easier to work with during stack construction. Two approaches
   that we are pursuing are to use a Gortex gasket with adhesive on the back
   and to make a v-shaped o-ring slot to capture the o-ring. Bill has
   ordered some sample gasket material and the JLab shop is making up a
   test o-ring. Stay tuned for developments.
 - We have to be out of the clean room by October 15. We will investigate
   way to bring the clean tent into room 126 and set it up there.

Wire Frame Update I

 - Wire frame P1 status:
    * Frame P1 has had its high voltage components fully sealed in
      Humiseal after drying out in the tent for several days. The currents
      were found to be acceptible on all busses. After the meeting, Bill
      machined in the pockets for the clearance of the rigid-flex assemblies
      on our single screwed up cathode. The mission was apparently successful.
      A voltage check will be done tomorrow and the frame will then be ready 
      for lamination to its associated cathode plane (this cathode still
      needs to have its rigid-flex assemblies soldered into place).
    * Due to handling, one of the preamp card connectors was broken and 
      will need to be replaced. Also one of the low voltage power connectors
      was broken off and several of the contact pads came with it.
 - Wire frame P3 status:
   * Frame P3 was set up such that its active area was setup in an
     Ar/CO2 gas environment, but the rest of the wire frame was left in
     air (just like the nominal configuration). Due to problems with getting
     a good gas seal, Snoop was squirted liberally on the boards. Snoop
     appears to be a pretty good conductor, and the currents on the HV
     busses rose to a level that the busses could not be even turned on
     to 10 V. After initial cleaning, the problems were slightly improved,
     but still horrible. In taking apart the gas setup to allow for access
     for cleaning, 3 wires were broken. Brian has replaced 2 wires and has
     one more to go.
   * We did get some encouraging news before the Snoop problem had taken
     hold. Beni measured the currents on some of our worst busses after
     gas flow had been going for about an hour and found them to be very
     low. However, full measurements and checkout was not finished before
     the Snoop wicked its way around the board and ended hope of doing
     sensible measurements.
   * At present we are still working to clean P3 and continue the gas
     tests. We are now moving to set up a system to clean the board more
     fully in alcohol and deionized water. We hope to have something set
     up in the next couple of days.
 - Wire frame P4 status:
   * P4 has current draw problems similar to P3, maybe a slight bit better.
     We only have current draw information at normal room humidity and we
     have let the board sit in the dry tent for a while. We will inspect
     the board and make current measurements when we get time. We will then
     do the gas test similar to what we did for frame P3.

Wire Frame Update II

 - In the full-scale prototype, there was a design problem with the circuit
   boards so the wire frames will be laminated to their neighboring cathode
   planes to give a gas seal. We will move forward with this work as we
   certify each wire frame as ready.
 - We need to have a dedicated meeting to design what the schedule for
   work on the Phase 3 wire winding will be. The highest priority for
   work scheduling will be doing work on upgrades of the wire winding
   facility. This includes a better strongback design, the design and
   construction of precision combs, modification to the wire winding
   table, and a list of other items with the system. Stay tuned, more
   to come.
 - The wire frame construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. We will include
   the information on the wire electroplating process shortly.
 - The parts of the jig to use for the wire electroplating operation
   are in our hands. They will be assembled as we find time.
 - Brian received the SEM scans for our wire from the W&M folks. He will
   generate a final QA report with photos and test results for all wires
   when he gets a chance.

Cathode Update

 - Progress updates:
     * The mechanical work on all cathodes is complete. The last thing
       to do is to connect the grounds on the individual panels on each
       cathode. A first attempt was made, but Fernando is not fully
       satisfied with the contact area and we will revisit it.
     * Lots of problems were found with the rigid-flex assemblies during
       electrical QA testing. Fernando was not pleased and took the boards
       back for additional checkout, repairs, and cleaning. He has since 
       certified a number of boards and made them ready for soldering to 
       the cathodes. The remainder of the rigid-flex assemblies should be 
       provided for soldering shortly.
     * Mark S. has not done any additional test soldering over the past
       week as we were awaiting a new heat sink from the shop. It is now
       in our possession and Mark S. will return to finalizing the soldering
       technique on test pieces. Certification of the soldering technique 
       still needs to take place before work with the real cathodes can begin.
 - The current draft of the cathode construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS.
 - Roger is pressing the Japanese company that has agreed to provide our
   board material. They have agreed to manufacture the material and it
   appears it will be delivered to JLab by the end of this year.
 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note. He indicated that he
   is back to work on this document and it should be done shortly.

Cosmic Ray Test Stand

 - Beni is continuing work to setup the cosmic ray telescope and test
   system in EEL 126. He is making progress on reducing the noise and is
   working to reproduce the data from the small-scale prototype that Simon
   had last year. The latest plots are in the FDC electronic logbook at:
 - We still have some mysterious behavior with the small-scale prototype
   response with wiggles in the X vs. Y distributions that need to be
   understood. Beni has reproduced this problem qualitatively and we
   discussed some possible causes, but more work needs to be done. Beni has
   just completed a pulser calibration run and his look at the small-scale
   prototype data including the new channel-to-channel gain matching (albeit
   with low statistics) looks very promising. Perhaps the wiggles are gone
   and can be blamed on the calibration.
 - Beni has started a bench test plan for the full-scale prototype and
   is presently working on it after incorporating feedback. The current
   information is contained on the FDC wiki page under the heading 
   "FDC test cosmic ray test stand".
 - The parts for the support/rotation of the full-scale prototype in the 
   cosmic ray stand will be assembled as we have time.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or 
   comments from the group.

Minutes prepared by Daniel. Send any comments or corrections along.